OpenStudy (anonymous):

jreis & limitless. Im trying to figure out how to solve the equation.. x/(x=12)=3/(x=3)

OpenStudy (anonymous):

(x+12)***

OpenStudy (anonymous):

\[\frac{x}{x+12}=\frac{3}{x+3}\]?

OpenStudy (anonymous):

satellite, yes.

OpenStudy (anonymous):

alright so cross multiply those mofos. you get x(x+3)=3(x+12) then distribute each side, x^2+3x=3x+36 now subtract 3x from each side, x^2=36 now square root both sides, x=-6 and 6

OpenStudy (anonymous):

\[\frac{x}{x+12}=\frac{3}{x+3}\] First, cancel the \(x+12\) and \(x+3\) by multiplying both sides by \((x+12)(x+3)\). This gives you: \[(x+3)x=3(x+12)\] Now you distribute the \(x\) and \(3\): \[x^2+3x=3x+36\] The \(3x\)'s cancel. So: \[x^2=36\] Take the squareroot of both sides: \[x=\pm\sqrt{36}=\pm 6\]

OpenStudy (anonymous):

^ what he said but less fancy

OpenStudy (anonymous):

\(\pm 6\) just means "plus or minus 6", so 6 and -6.

OpenStudy (anonymous):

K, but... cross multiply. like, what do you mean by that? and haha Jreis.lol

OpenStudy (anonymous):

Cross multiply just means multiplying both sides (of the equation) by \((x+12)(x+3)\).

OpenStudy (anonymous):

so its basically the denominators?

OpenStudy (anonymous):

Yes. Precisely! :D

OpenStudy (anonymous):

so itd be x times (x+12)(x+3)

OpenStudy (anonymous):

Well, no. The \((x+12)\) goes away: \[(x+12)(x+3)\frac{x}{x+12}=\frac{(x+12)(x+3)x}{x+12}=\frac{(x+12)}{x+12}(x+3)x=(x+3)x\] \(\frac{x+12}{x+12}\) is just 1.

OpenStudy (anonymous):

cross multiply means you take the denominator from left side and move it to the numerator of the right, and vice versa. take the denominator from one side and bring to the other, so (x+12) goes to right, (x+3) goes to left